OpenStudy (anonymous):

how to solve this equation: X(1+2X)=(2X-1)(X-2)

OpenStudy (anonymous):

Distribute all out: 2x² + x = ... Can you continue?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

is it x+2x^2=2x^2-4x-x+2 ?

OpenStudy (anonymous):

okay so what do i do next

OpenStudy (anonymous):

okay so i got x=1/3 is that right?

OpenStudy (anonymous):

Absolutely correct :D

OpenStudy (anonymous):

thank you!! :)